OpenStudy (anonymous):

2x-2=x+64 help me solve it plzz ^_*

OpenStudy (gitface):

\[2x-2=x+64\] Begin by grouping your variable onto one side only. This can be accomplished by either subtracting 2x from both sides, or subtracting only 1x from both sides. In our case, it would be easier to subtract 1x from both sides. \[2x-x-2=x-x+64\]\[x-2=64\] Now that your variable is only on one side, make it the only term on that side. This is done by subtracting 2 from each side. \[x-2+2=64+2\]\[x=66\] Now the equation is such that your variable is alone on one side and is a single variable. The equation cannot be simplified any further, so you are finished.